OpenStudy (anonymous):

If y varies directly as x, what is the value of y in these ordered pairs? (1, 1) and (17, y)

OpenStudy (zzr0ck3r):

y=kx 1=1*k so k =1 so y = x so (17,17)

OpenStudy (adilalvi):

17

OpenStudy (zzr0ck3r):

y=kx (y varies with x) and we use the point (1,1) to find k plug in y=1 and x = 1 y = kx becomes 1 = k*1 and so we know that k = 1 so y=x now the last pair is (17,y) and we know y = x = 17 so y = 17 thus (17,17)